On 1/1/2011, ABC Company purchased a piece of equipment costing $65,000. The equipment is expected to have a useful life equal to 5 years and a salvage value equal to $5,000.

Calculate the first two years’ depreciation expense using the following methods respectively.

1) the straight-line method
2011:
2012:

2) the double-declining method and
2011:
2012:

3) the sum-of-years’-digits’ method.
2011:
2012:
Answer #1

Correct Answer:

Requirement 1:

Year

Depreciation expense

2011

$12,000.00

2012

$12,000.00

Working:

Straight Line Depreciation

A

Cost

$ 65,000.00

B

Residual Value

$ 5,000.00

C=A - B

Depreciable base

$ 60,000.00

D

Life [in years]

5

E=C/D

Annual SLM depreciation

$ 12,000.00

Requirement 2:

Year

Depreciation expense

2011

$ 26,000.00

2012

$ 15,600.00

Working:

Double Declining Method

A

Cost

$ 65,000.00

B

Residual Value

$ 5,000.00

C=A - B

Depreciable base

$ 60,000.00

D

Life [in years]

5

E=C/D

Annual SLM depreciation

$ 12,000.00

F=E/C

SLM Rate

20.00%

G=F x 2

DDB Rate

40.00%

Year

Beginning Book Value

Depreciation rate

Depreciation expense

Ending Book Value

1

$ 65,000.00

40.00%

$ 26,000.00

$ 39,000.00

2

$ 39,000.00

40.00%

$ 15,600.00

$ 23,400.00

Requirement 3:

Year                                                              Depreciation expense

2011

$ 20,000

2012

$ 16,000

Working:

Sum of year's digit

Cost

$ 65,000.00

Residual Value

$ 5,000.00

Depreciable base

$ 86,520.00

sum of years (1+2+3+4+5) = 15

15

                                  2011

$ 20,000

Depreciable base * ( 5/15)

                              2,012

$16,000

Depreciable base * (4/15)

End of answer.

Please give a thumbs-up, it will be highly appreciated.

Thanks.